undeadknight26 (undeadknight26):

Derive the equation of the parabola with a focus at (0, -4) and a directrix of y = 4.

undeadknight26 (undeadknight26):

I got f(x) = 16x^2

OpenStudy (anonymous):

Still using that picture, eh?

OpenStudy (anonymous):

Pretty embarrassing tbh.

undeadknight26 (undeadknight26):

Ayeeeee XD

undeadknight26 (undeadknight26):

Its fine by me lel

OpenStudy (haseeb96):

for equation of Parabola the focus is given that is (0,-4) and the directrix y=4

OpenStudy (haseeb96):

here a =-4 and this parabola is along y axis so the equation would be in this form x=4ay now just put a =-4 in the above equation and get the answer

OpenStudy (haseeb96):

*x^2 =4ay

OpenStudy (haseeb96):

x^2 = 4(-4) y x^2 =-16 y is the equation of parabola

undeadknight26 (undeadknight26):

sooo: x^2 = 4(-4)y? x^2 = -16y? should i now invert it so its: y = -16x^2?

OpenStudy (haseeb96):

no when focus =(a,0) is on this condition then we will use this equation parabola that is y^2=4ax and when focus =(0,a) is on this condition then we will use this equation parabola that is x^2=4ay okay i think u got it

undeadknight26 (undeadknight26):

Can you help me wif one more mate?